OpenStudy (anonymous):

the average of fraction is 3/8. one of the fractions is 1/2. find the other fraction?

OpenStudy (anonymous):

3/8 *2 -1/2=1/4

OpenStudy (ash2326):

let the other fraction be x given the average of the two fractions is 3/8 (1/2+x)/2=3/8 1/2+x=6/8 1/2+x=3/4 subtract 1/2 from both sides x=3/4-1/2 x=3/4=2/4 x=1/4 so the other fraction is 1/4

OpenStudy (anonymous):

why did you divide by two?

OpenStudy (ash2326):

average of two numbers a and b is (a+b)/2 similarly average of three is calculated by (a+b+c)/3